FORMATION OF QUASARS AT HIGH REDSHIFT

Neal Katz^{1}^{,}^{2}, Thomas Quinn^{2,}^{3}, Edmund Bertschinger^{1}, and James M. Gelb^{1,}^{4}

Received:

ABSTRACT Analytic arguments have suggested that the existence of high redshift quasars challenge hier-

archical galaxy formation theories where the majority of galaxies form late. We address this issue using a series of numerical simulations of the biased cold dark matter model with and without gas dynamics. Potential quasar sites are identified with massive density peaks collapsing to form massive self-gravitating cold gas clouds. The Press-Schechter formalism correctly predicts the number of high mass dark halos using a Gaussian filter with δ_{c} = 1.61. However, very efficient cooling and angular momentum transfer to the dark halo allows the gas to become much denser than naive predictions. We find there are > 10^{3} times the required number of potential quasar sites at z = 4.5; at z = 8 there is already at least one site in our simulation volume of 10^{6} Mpc^{3}. Thus, sites for quasars could easily form at significantly higher redshifts than observed quasars, even in models where most galaxy formation occurs late. Subject headings: galaxies: formation — quasars

1 Department of Physics, MIT Room 6-207, Cambridge, MA 02139

2 Department of Astronomy FM-20, University of Washington, Seattle, WA 98195 (current address)

3 Theoretical Physics, Keble Road, Oxford OX1 3NP

4 NASA/Fermilab Astrophysics Center, P.O. Box 500, Batavia, IL 60510